93?35?175?26?128?16=1 in the place of question mark eithe + or - 'm' be the no.of + sign & 'n' be the no.of - sign so m-n=?
Answer by Edwin McCravy(20056)

The only way to get 1 is

93-35-175-26+128+16 = 1

There are 2 + signs and 3 - signs 
So m=2 and n=3, so m-n = 2-3 = -1

Edwin